# Intention Knowledge Graph Construction for User Intention Relation Modeling

### v1

@@ 5 Extrinsic Evaluation
Since our datasets are from the real-world Amazon M2 recommendation datasets, it is natural and necessary to validate the effectiveness of our work on the corresponding downstream tasks to show its usefulness. With this section, we demonstrate that our KG can provide extra information gain for session recommendation, which is also one of the major goals of the current KG.
Datasets Metric FPMC GRU4Rec BERT4Rec SASRec SASRecF CORE SR-GNN GCE-GNN DIF-SR FEARec RIGRec
M2 (UK) Recall@5 0.2523 0.2792 0.1899 0.3075 0.2957 0.2990 0.2928 0.3130 0.3128 0.3088 $\textbf{0.3342}^{*}$
Recall@10 0.3121 0.3469 0.2641 0.3964 0.3713 0.3949 0.3678 0.4001 0.3990 0.3941 $\textbf{0.4229}^{*}$
Recall@20 0.3696 0.4108 0.3349 0.4723 0.4406 0.4768 0.4381 0.4726 0.4739 0.4691 $\textbf{0.5003}^{*}$
Recall@50 0.4389 0.4865 0.4197 0.5621 0.5245 0.5697 0.5171 0.5542 0.5598 0.5552 $\textbf{0.5863}^{*}$
Recall@100 0.4841 0.5346 0.4744 0.6159 0.5771 0.6223 0.5662 0.6032 0.6171 0.6100 $\textbf{0.6398}^{*}$
NDCG@5 0.1933 0.2118 0.1260 0.2121 0.2208 0.1673 0.2195 0.2214 0.2171 0.2138 0.2214
NDCG@10 0.2126 0.2327 0.1501 0.2406 0.2432 0.1985 0.2418 0.2441 0.2451 0.2415 $\textbf{0.2503}^{*}$
NDCG@20 0.2272 0.2499 0.1682 0.2598 0.2634 0.2193 0.2616 0.2626 0.2641 0.2605 $\textbf{0.2703}^{*}$
NDCG@50 0.2411 0.2648 0.1848 0.2679 0.2801 0.2379 0.2784 0.2807 0.2812 0.2777 $\textbf{0.2877}^{*}$
NDCG@100 0.2484 0.2718 0.1937 0.2862 0.2887 0.2472 0.2865 0.2871 0.2891 0.2866 $\textbf{0.2957}^{*}$
@@
5.1 Data Preparation
The whole English subset of the M2 dataset with our generated knowledge graph is utilized in this step for the session recommendation task. We split all sessions into train, validation, and test sets at a ratio of 8:1:1. Since the original dataset is prepossessed before release, we do not employ additional filtering operations to avoid the risk of session loss or false connections among items.
Since all of the current session recommendation paradigms are difficult to integrate a million-level commonsense knowledge graph as we constructed in an end-to-end manner, we build upon an item relation graph based on it with meta-path methods to describe product relations. Specifically, we first gather all 1-hop session pairs where concepts or temporal relations directly connect their intentions. To alleviate redundant connection and noise, we only keep session pairs if (1) they have no less than six distinct meta paths passed through commonsense relation nodes or (2) one can reach the other from all of its 1-hop concept nodes. Then, we construct a weight item graph $G=(V,E)$, where $V$ is the node-set, $E$ is the edge set generated by linked items in all sampled session pairs, and the co-occurrence frequency is as weight. Here, we focus on evaluating the effectiveness of relations derived from our session-intention KG, leaving the exploration of optimal sub-graph sampling strategy to be explored in future work.
5.2 Graph Representations for Session Recommendation
Given the relation $G$, one can obtain item representations with various graph representation methods. We adopt the simple yet effective graph convolution operation to learn informative item representations, which is formulated as
$\bm{E}^{l+1}=\bm{A}\bm{E}^{l},$
where $\bm{A}$ is the adjacency matrix of $G$, $\bm{E}\in\mathbb{R}^{N\times d}$ is the $d$-dimensional embedding dictionary of all items. Canonical methods are utilized during graph representation learning, including light-weight convolution and sum pooling. After $L$-th convolution, the knowledge graph-based representations $\bm{E}^{L}$ can be directly utilized in the recommendation tasks for enhancing performance.
To achieve end-to-end recommendation, we designed a Relational Intention Knowledge Graph-based recommender named RIGRec, which seamlessly integrates the graph representations learning module into the session recommendation framework. Concretely, the widely used attention-based method SASRec is adopted as our session encoder, which aggregates the learned knowledge graph-based item representations within each session for user preference estimation.
@@ 5.2 Graph Representations for Session Recommendation
5.3 Baselines and Evaluation Metrics
@@
We compare our model with following ten representative and state-of-the-art methods, covering (1) the classical method FPMC Rendle et al. (2010), (2) the RNN-based method GRU4Rec Hidasi et al. (2016), (3) the predominant attention-based methods including BERT4Rec Sun et al. (2019), SASRec Kang and McAuley (2018), CORE Hou et al. (2022) and FEARec Du et al. (2023), (4) graph-based methods including SR-GNN Wu et al. (2019b) and GCE-GNN Wang et al. (2020), (5) side information fusion methods including SASRecF Kang and McAuley (2018) and DIF-SR Xie et al. (2022). We exclude some state-of-the-art methods like FAPAT Liu et al. (2023b) due to the need for massive support resources or exponential computation complexity.
@@
We employ two standard evaluation metrics in the field of recommender systems, including Recall at a cutoff top $k$ (Recall@$k$) and Normalized Discounted Cumulative Gain at a cutoff top $k$ (NDCG@$k$). We rank the ground-truth item alongside all candidates to ensure an unbiased evaluation rather than adopting the negative sampling strategy. We report the averaged metrics over 5 runs with the commonly utilized $k\in\{5,10,20,50,100\}$. The implementation details are in the Appendix A.5.
5.4 Performance Comparison
Table 10 presents the recommendation performance of all models on the M2 dataset. An unpaired T-test with a p-value of 0.05 is conducted to prove the improvement is statistically significant. From these results, We have the following observations.
Our method consistently surpasses all baselines by a considerable margin regarding most metrics. Performance improvement should be attributed to the major characteristic of our model, i.e., refining useful information from the intention knowledge graph to enhance recommendations. More persuasive evidence is the performance discrepancy between our RIGRec and SASRec, which serves as the backbone of our method, further verifying the effectiveness of our intention knowledge graph.
Graph-based methods SR-GNN and GCE-GNN present inferior effectiveness compared to our model, which is not a surprise since they only capture transition relation among items, quickly leading to unsatisfactory modeling. In contrast, we endow item representations with adequate complex dependencies acquired from our intention knowledge graph and thus achieve improvement. This justifies the necessity of enriching recommendation tasks with amply modeling user intentions.
Compared to most baselines, DIF-SR integrates side information for recommendation with an elaborately designed structure to achieve competitive performance. In contrast, another side information fusion method, SASRecF, exhibits performance deterioration compared to SASRec in terms of Recall@$k$. Such an observation is consistent with previous works Xie et al. (2022); Liu et al. (2021), suggesting that inappropriate information fusion strategies may degenerate the model efficiency. Besides, noise and the inevitable increase of defect value with the expansion of the data scale also limit their performance. On the contrary, we resort to inferring user intentions with the aid of LLM, alleviating the aforementioned problems by reducing ineffectual information from raw features of items.
5.5 Ablation Study
To research the effectiveness of information provided by conceptualization and concepts relation of intentions, we conduct ablation studies by comparing our RIGRec with three variants generated by removing different types of edges as follows: (1) w/o all: This variant removes the whole item graph, equal to SASRec. (2) w/o concept: This variant removes edges acquired by the meta path through concepts. (3) w/o commonsense relation: This variant removes edges acquired by meta path through commonsense relation nodes.
Figure 2 presents the results of the ablation studies; we interpret the results with the following discoveries. First, our constructed intention knowledge graph is of vital use to produce helpful item representations for downstream recommendation tasks. As we can see, the performance of our model exceeds that of SASRec by a large margin. Second, both types of information contained in intention conceptualization and concepts relation prove efficient. It can be confirmed that the two variants solely with corresponding edges can outperform w/o all variants. Third, distinct relations provide dissimilar contributions to each metric, w/o commonsense relation presenting high Recall metrics. At the same time, w/o conceptualization performs better in NDCG metrics, which reaffirms the significance of constructing multi-type nodes and relations in our intention knowledge graph.
@@ 6 Conclusion
We present IGC-RC, a framework for automatically constructing intentional commonsense knowledge graphs from user behaviors. Using this framework, we built the Relational Intention Knowledge Graph (RIG) and validated its quality through extensive evaluations. Results show that RIG significantly enhances the performance of state-of-the-art session recommendation models.
@@ Appendix C Annotation Questions
Here, we give three examples of annotation questions from our questionnaire for the Amazon Mechanical Turk. They are for the session intentions quality annotation, the intention relation classification annotation, and the intention conceptualization classification.

### v2

@@ 5 Extrinsic Evaluation
To demonstrate the practical utility of our relational intention knowledge graph, we evaluate its effectiveness in enhancing session-based recommendation, a key application domain for e-commerce platforms. Using the Amazon M2 dataset, we show how the rich semantic relationships captured in our knowledge graph can improve recommendation quality compared to state-of-the-art methods.
Datasets Metric FPMC GRU4Rec BERT4Rec SASRec SASRecF CORE SR-GNN GCE-GNN DIF-SR FEARec DGNN ISRec Satori KA-MemNN RIGRec
M2 (UK) Recall@5 0.2523 0.2792 0.1899 0.3075 0.2957 0.2990 0.2928 0.3130 0.3128 0.3088 0.3021 0.3073 0.2973 0.2932 $\textbf{0.3342}^{*}$
Recall@10 0.3121 0.3469 0.2641 0.3964 0.3713 0.3949 0.3678 0.4001 0.3990 0.3941 0.3882 0.3981 0.3821 0.3781 $\textbf{0.4229}^{*}$
Recall@20 0.3696 0.4108 0.3349 0.4723 0.4406 0.4768 0.4381 0.4726 0.4739 0.4691 0.4563 0.4754 0.4436 0.4419 $\textbf{0.5003}^{*}$
Recall@50 0.4389 0.4865 0.4197 0.5621 0.5245 0.5697 0.5171 0.5542 0.5598 0.5552 0.5584 0.5676 0.5348 0.5295 $\textbf{0.5863}^{*}$
Recall@100 0.4841 0.5346 0.4744 0.6159 0.5771 0.6223 0.5662 0.6032 0.6171 0.6100 0.6072 0.6201 0.5931 0.5847 $\textbf{0.6398}^{*}$
NDCG@5 0.1933 0.2118 0.1260 0.2121 0.2208 0.1673 0.2195 0.2214 0.2171 0.2138 0.2201 0.2207 0.2189 0.2163 0.2214
NDCG@10 0.2126 0.2327 0.1501 0.2406 0.2432 0.1985 0.2418 0.2441 0.2451 0.2415 0.2431 0.2438 0.2425 0.2386 $\textbf{0.2503}^{*}$
NDCG@20 0.2272 0.2499 0.1682 0.2598 0.2634 0.2193 0.2616 0.2626 0.2641 0.2605 0.2637 0.2633 0.2623 0.2591 $\textbf{0.2703}^{*}$
NDCG@50 0.2411 0.2648 0.1848 0.2679 0.2801 0.2379 0.2784 0.2807 0.2812 0.2777 0.2797 0.2795 0.2793 0.2737 $\textbf{0.2877}^{*}$
NDCG@100 0.2484 0.2718 0.1937 0.2862 0.2887 0.2472 0.2865 0.2871 0.2891 0.2866 0.2867 0.2881 0.2864 0.2843 $\textbf{0.2957}^{*}$
@@
Data Preparation
We use the complete English subset of the Amazon M2 dataset integrated with our constructed knowledge graph. Following standard practice, we partition all sessions into training, validation, and test sets at an 8:1:1 ratio. To preserve the integrity of the dataset, we use the original preprocessing without filtering, thus avoiding potential session loss or artificial item connections.
The integration of a large-scale commonsense knowledge graph (containing over 351 million edges) with conventional recommendation systems presents significant challenges. To address this, we develop a meta-path approach to extract actionable item relationships from our knowledge graph. First, we identify session pairs connected through intention-level relationships (either conceptual or temporal). To ensure relationship quality, we retain only session pairs that either (1) share at least six distinct meta-paths through commonsense relation nodes or (2) demonstrate complete reachability from one session’s concept nodes to the other’s. From these high-quality session pairs, we construct a weighted item graph $G=(V,E)$, where nodes represent products and edges represent their co-occurrence relationships within semantically connected sessions. Edge weights correspond to co-occurrence frequency, providing a measure of relationship strength. This approach effectively distills our comprehensive knowledge graph into a focused item relationship network tailored for recommendation tasks.
RIGRec: Intention-Enhanced Recommendation Model
We develop RIGRec, a novel recommendation model that leverages the rich semantic relationships captured in our intention knowledge graph. The model employs graph convolution to derive informative item representations:$\bm{E}^{l+1}=\bm{A}\bm{E}^{l},$ where $\bm{A}$ represents the adjacency matrix of graph $G$, and $\bm{E}\in\mathbb{R}^{N\times d}$ stores the $d$-dimensional embeddings for all $N$ items. We implement lightweight convolution operations and sum pooling to enhance computational efficiency. After $L$ convolution layers, the resulting representations $\bm{E}^{L}$ capture both item characteristics and their intention-based relationships. For session modeling, we adopt SASRec’s self-attention architecture, which effectively aggregates item representations within a session to model user preferences. This creates an end-to-end recommendation system that seamlessly integrates intention-level reasoning with sequential pattern recognition, enabling more contextually appropriate recommendations.
@@
Baselines and Evaluation Metrics
@@
We compare our model with following ten representative and state-of-the-art methods, covering (1) the classical method FPMC Rendle et al. (2010), (2) the RNN-based method GRU4Rec Hidasi et al. (2016), (3) the predominant attention-based methods including BERT4Rec Sun et al. (2019), SASRec Kang and McAuley (2018), CORE Hou et al. (2022) and FEARec Du et al. (2023), (4) graph-based methods including SR-GNN Wu et al. (2019b) and GCE-GNN Wang et al. (2020), (5) side information fusion methods including SASRecF Kang and McAuley (2018) and DIF-SR Xie et al. (2022). We exclude some state-of-the-art methods like FAPAT Liu et al. (2023b) due to the need for massive support resources or exponential computation complexity. Additionally, we compare our approach with state-of-the-art intention-aware recommendation models: DGNN Li et al. (2023), ISRec Li et al. (2021b), Satori Chen et al. (2022), and KA-MemNN Zhu et al. (2021).
@@
We employ two standard metrics in the field of recommender systems, including Recall at a cutoff top $k$ (Recall@$k$) and Normalized Discounted Cumulative Gain at a cutoff top $k$ (NDCG@$k$). We rank the ground-truth item alongside all candidates to ensure an unbiased evaluation rather than adopting the negative sampling strategy. We report the averaged metrics over 5 runs with the commonly utilized $k\in\{5,10,20,50,100\}$. To ensure unbiased evaluation, we rank the ground-truth item alongside all candidate items rather than using negative sampling. Results represent averages across five independent runs, with statistical significance determined through unpaired t-tests (p < 0.05). The implementation details are in the Appendix A.5.
Performance Comparison
As shown in Table 10, RIGRec consistently outperforms all baseline methods across nearly all evaluation metrics, with statistically significant improvements over the strongest competitors. These results demonstrate the substantial benefits of incorporating intention-level relationships.
The performance gap between RIGRec and SASRec—which serves as RIGRec’s backbone architecture—is particularly notable. This difference directly quantifies the value added by our intention knowledge graph, confirming that modeling user intentions and their relationships provides critical information beyond what can be captured by sequential patterns alone.
Graph-based methods like SR-GNN and GCE-GNN, while theoretically capable of modeling complex item relationships, show inferior performance compared to our approach. This stems from their focus on direct item transitions without consideration of the underlying user intentions driving these transitions. By explicitly modeling intention-level relationships, RIGRec captures deeper semantic connections between items, resulting in more contextually appropriate recommendations.
The mixed performance of side information fusion methods (DIF-SR and SASRecF) highlights the challenges of effectively integrating auxiliary information into recommendation systems. While DIF-SR achieves competitive results through careful information fusion, SASRecF shows performance degradation compared to the base SASRec model in terms of Recall@k. This aligns with findings that inappropriate information integration strategies can degrade model effectiveness. Our approach, by contrast, leverages large language models to extract high-quality intention information, filtering out noise and irrelevant features that might otherwise compromise recommendation quality.
Ablation Study
To isolate the contributions of different components in our knowledge graph, we conduct ablation studies comparing RIGRec against three variants: (1) w/o all: Removes the entire item graph derived from our knowledge graph (equivalent to standard SASRec). (2) w/o concept: Removes edges derived from concept-mediated paths. (3) w/o commonsense relation: Removes edges derived from commonsense relation paths.
The ablation study reveals three insights: (1) The intention knowledge graph substantially improves recommendation quality beyond what sequential patterns alone can achieve; (2) Both conceptualization and commonsense relations provide complementary signals that enhance performance; and (3) These relation types contribute differently to evaluation metrics—conceptualization improves recall by identifying broadly relevant items, while commonsense relations enhance NDCG by capturing fine-grained semantic connections for better ranking. These findings confirm that modeling diverse intention relationships leads to more contextually appropriate recommendations by providing a more comprehensive understanding of user behavior.
@@ 6 Conclusion
We introduce IGC-RC, a framework for automatically constructing knowledge graphs that model relationships between user intentions in e-commerce contexts. Our Relational Intention Graph (RIG) captures temporal, causal, and conceptual connections between user goals, demonstrating exceptional quality in both intrinsic evaluations and recommendation tasks. By bridging the gap between observable behaviors and underlying intentions, RIG enables more accurate prediction of user needs and significantly enhances recommendation performance. This advancement represents an important step toward more human-like understanding in intelligent e-commerce systems.
@@ A.5 Implementation Details of Session Recommendation
N-gram FolkScope RIG w/ConceptNet Rel
2-gram 0.0307 0.5274 0.3759
3-gram 0.0480 0.7931 0.5806
4-gram 0.0648 0.9100 0.6997
5-gram 0.0837 0.9623 0.7751
6-gram 0.1046 0.9833 0.8264
Table 11: N-gram diversity scores of intentions extracted from FolkScope and RIG. The results validate that RIG generates more diverse intentions by removing ConceptNet relation constraints.
@@ Appendix B Large Language Model Generation Prompts
Session Intention Generation Below is a user’s chronological record list: [SESSION] Explain the basic intentions of this user exactly. Output several different intentions one by one to answer the following question: Users buy these items because they want to: intention 1: {a simple verb phrase within 10 words} intention 2: {a simple verb phrase within 10 words} …
@@
Abstract Intention Generator I will give you an INTENTION. You need to give several phrases containing 1-3 words for the ABSTRACT INTENTION of this INTENTION. You must return your answer in the following format: phrases1,phrases2,phrases3,…., which means you can’t return anything other than answers. These abstract intention words should fulfill the following requirements: 1. The ABSTRACT INTENTION phrases can well represent the INTENTION. 2. The ABSTRACT INTENTION phrases don’t have a lot of less relevant word meanings. For example, “spring” is not a good abstract intention word because it can represent both a coiled metal device and the season of the year. 3. The ABSTRACT INTENTION phrases of the same INTENTION cannot be semantically similar to each other. For example, health and wellness are two close synonyms, so they can’t be together. INTENTION: Moisturize dry skin while enjoying a special effect bath. Your answer: hydration, skincare INTENTION: Create a festive atmosphere for a Christmas party. Your answer: party planning, celebration, decorations, holiday spirit INTENTION: [INTENTION]. Your answer:
@@ Appendix C Annotation Questions
Here, we give three examples of annotation questions from our questionnaire for the Amazon Mechanical Turk. They are for the intentions quality annotation, intention relation classification annotation, and intention conceptualization classification.
Model R@5 R@10 R@20 R@50 R@100 N@5 N@10 N@20 N@50 N@100
GRU4Rec 0.2792 0.3469 0.4108 0.4865 0.5346 0.2118 0.2327 0.2499 0.2648 0.2718
RIG-GRU4Rec 0.2923 0.3852 0.4665 0.5562 0.6028 0.2126 0.2367 0.2474 0.2693 0.2746
SASRec 0.3075 0.3964 0.4723 0.5621 0.6159 0.2121 0.2406 0.2598 0.2679 0.2862
RIGRec 0.3342 0.4229 0.5003 0.5863 0.6398 0.2214 0.2503 0.2703 0.2877 0.2957
Table 13: Performance comparison of different backbone architectures with and without our knowledge graph.
Appendix D Evaluating Diversity of Intention
As shown in Figure 11, we further compare the diversity of the generated intentions by using the n-gram diversity. It is defined as the ratio of the unique n-gram counts to all n-gram counts:
$\text{Diversity}(D,n)=\frac{\#\text{ unique }n\text{-grams in }D^{\oplus}}{\#\text{ n-grams in }D^{\oplus}}$ (1)
Where $D^{\oplus}$ denotes the dataset $D$ concatenated into a single string. We use six as the maximum n-gram length. This method captures repeated sequences in addition to single-token diversity.
We measured the diversity of the corpus formed by the intentions extracted from ForkScope and RIG. For the intention generation with conceptnet relation, we randomly sample 100 intentions derived from each of 17 conceptnet relations, totaling 1700 intentions. For RIG and Folkscope, we sample 1700 intentions from each of them to compute the n-gram diversity. The results demonstrate that RIG’s N-gram diversity of intentions is significantly higher than ForkScope’s. These findings validate our claim that removing ConceptNet relation constraints and using a better generation model generates more diverse intentions.
Appendix E Analysis of Different Backbones for RIGRec
The choice of backbone architecture is an important consideration when implementing our approach. In our main experiments, we primarily utilized SASRec as the backbone for RIGRec due to its simplicity, robustness, and widespread adoption as a baseline in session-based recommendation research. This choice provides clear interpretability of the improvements brought by our intention knowledge graph.
However, to demonstrate the generalizability of our approach, we conducted additional experiments using GRU4Rec as an alternative backbone. Table 13 presents the performance comparison between the original backbones (GRU4Rec and SASRec) and their RIG-enhanced versions (RIG-GRU4Rec and RIGRec).
The results demonstrate that our approach consistently improves performance regardless of the backbone architecture. When applied to GRU4Rec, our intention knowledge graph enhances performance across all metrics, with particularly notable improvements in Recall@20 (13.6% improvement) and Recall@50 (14.3% improvement). This confirms that the benefits of our relational intention knowledge graph are not limited to a specific recommendation architecture.
Nevertheless, the RIGRec model (based on SASRec) still achieves the best overall performance. This can be attributed to SASRec’s inherent advantages in capturing long-range dependencies through its self-attention mechanism, which may better complement the high-level intention information provided by our knowledge graph. The results suggest that while our approach generalizes well to different backbones, the choice of backbone can still influence the absolute performance levels achieved.
Appendix F Case Studies
To illustrate the practical effectiveness of our RIG approach, we present three representative case studies in Table 14. These examples highlight how our model leverages intention relationships to generate more contextually relevant recommendations compared to baseline methods.
Session History Generated Intentions Baseline Recommendations (SASRec) Our Model Recommendations (RIGRec)
Wireless Gaming Mouse; RGB Gaming Keyboard; Gaming Headset with Microphone Set up a complete gaming peripherals system; Enhance gaming experience with RGB accessories; Improve communication during multiplayer games Ergonomic Gaming Chair; Gaming Mouse Pad; USB Hub; Mechanical Gaming Keyboard; RGB Gaming Mouse Gaming Mouse Pad with RGB; Gaming Desk with Cable Management; Dual Monitor Stand; RGB LED Strip for Desk; Gaming Controller
Baby Diapers Size 1 (240 Count); Baby Wipes Unscented; Diaper Rash Cream; Baby Bottle Sterilizer Prepare essential baby care items for a newborn; Maintain baby hygiene; Prevent and treat diaper rash; Ensure safe feeding equipment Baby Powder; Baby Wipes Sensitive; Baby Lotion; Baby Shampoo; Diaper Bag Baby Bottles Anti-Colic; Bottle Drying Rack; Changing Pad; Baby Bottle Brush Set; Bottle Warmer
Pumpkin Carving Kit; Halloween String Lights; Artificial Spider Web; Halloween Doorbell with Spooky Sounds Decorate home for Halloween; Create a spooky atmosphere; Prepare for trick-or-treaters; Host a Halloween party Halloween Window Decorations; Halloween Candy Bowl; Halloween Mask; Halloween Costume; Fog Machine Halloween Door Wreath; Motion-Activated Halloween Props; Outdoor Halloween Projector Lights; Halloween Themed Serving Tray; Halloween Party Snack Bowls
Table 14: Case Study: Example Sessions and Recommendations
F.1 Gaming Setup Completion
In the first case, the user has purchased basic gaming peripherals (mouse, keyboard, and headset). The baseline model recommends similar gaming peripherals and accessories without considering the user’s broader intention. In contrast, our RIGRec model identifies the user’s goal of setting up a complete gaming system and recommends complementary products that enhance the overall gaming environment, such as a desk with cable management and RGB lighting solutions. This demonstrates how our model captures the conceptual relationships between products through the user’s intentions.
F.2 Baby Care Essentials
The second case shows a user purchasing basic newborn care items. While the baseline recommends additional hygiene products that are similar to those already in the cart, our model recognizes the broader intention of preparing for a newborn’s feeding needs. By identifying the connection between diaper changing and feeding essentials as part of comprehensive baby care, RIGRec suggests complementary products like anti-colic bottles and a bottle warmer that the baseline misses entirely.
F.3 Halloween Preparation
In the third case, the user is collecting Halloween decorations. The baseline model focuses on recommending additional decorative items and costumes. Our model, however, identifies the potential intention of hosting a Halloween party and recommends party-specific items like themed serving trays and snack bowls. This demonstrates how RIGRec’s understanding of intention relationships (decorating for Halloween → hosting a Halloween party) enables it to anticipate future user needs that aren’t explicitly indicated in the current session.
F.4 Analysis of Intention Relations
Table 15 shows examples of how our model captures key intention relationships that help generate better recommendations. These intention pairs demonstrate the temporal and causal connections our knowledge graph identifies.
Initial Intention Related Intention Relation Type
Set up a gaming peripherals system Create an immersive gaming environment Causality
Decorate home for Halloween Host a Halloween party Temporal (Before→After)
Ensure safe feeding equipment Prepare bottles for infant feeding Synchronous
Upgrade PC components Enhance gaming performance Causality
Purchase cooking utensils Prepare homemade meals Temporal (Before→After)
Table 15: Examples of Captured Intention Relations in RIG
These cases illustrate how our intention knowledge graph helps bridge the gap between observed behaviors and underlying user goals, resulting in recommendations that better address the user’s complete needs rather than simply suggesting similar products. The ability to identify conceptual relationships between different intentions enables our model to make more contextually appropriate and diverse recommendations.